solve by elimination method 0.3x -0.2y = 4 and 0.4x+0.3y=5/19 system of equations
Answer by checkley77(12844)

0.3x-0.2y=4 multiply by 3
0.4x+0.3y=5/19 myultiply by 2 & add.
.9x-.6y=12
.8x+.6y=10/19 add.
------------------
1.7x=12+10/19
1.7x=(12*19+10)/19
1.7x=(228+10)/19
1.7x=238/19
x=(238/19)/1.7
x=238/19*1.7
x=238/32.3
x=14/1.9 ans.
.3*7/.95-.2y=4
2.1/.95-.2y=4
-.2y=4-2.1/.95
-.2y=(4*.95-2.1)/.95
-.2y=(3.8-2.1)/.95
-.2y=1.7/.95
y=(1.7/.95)/-.2
y=1.7/.95(1/-.2)
y=-1.7/.19
y=-17/1.9 ans.
Proof:
.3(14/1.9)-.2(-17/1.9)=4
4.2/1.9+3.4/1.9=4
7.6/1.9=4
4=4